# Weighing the Axion with Muon Haloscopy

### v38

@@ Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the standard model of particle physics. We argue that axions form the dark matter in the universe and that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument gives a percent level prediction for the mass of the axion that can be tested by measuring the mid-infrared electrodynamic response of a wide range of polar materials. A broad resonance in the publicly available Raman spectrum from quartz is consistent with the creation of axions with the predicted mass. The difference in the frequency of a pair of absorption lines in the published infrared spectrum of polystyrene gives an estimate of the mass of the axion to 44 parts per million precision that is consistent with the predicted value.
@@
Muon Haloscopy—Axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=(m_{A}c)^{2}\ll(m_{\mu}c)^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [12] changes the photon’s Green’s function [13, 14]

### v39

@@ Abstract
Recent measurements of muon spin precession and an emerging consensus for the muon magnetic dipole moment confirm a long-standing tension within the standard model of particle physics. We argue that axions form the dark matter in the universe and that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument gives a percent level prediction for the mass of the axion that can be tested by measuring the mid-infrared electrodynamic response of a wide range of polar materials.
@@
Muon Haloscopy—Axions in the local dark matter halo shift muon spin precession [11]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=(m_{A}c)^{2}\ll(m_{\mu}c)^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [12] changes the photon’s Green’s function [13, 14]